# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

### v2

@@ D.2 Covariate Shift Experiments
CLIP RN50 0.94 0.91 0.17 0.92 0.89 0.23 0.93 0.90 0.27 0.92 0.84 0.19
RN101 0.96 0.94 0.15 0.92 0.89 0.25 0.87 0.83 0.39 0.91 0.84 0.22
RN50x4 0.98 0.98 0.07 0.94 0.93 0.19 0.87 0.83 0.43 0.83 0.75 0.34
RN50x16 0.97 0.96 0.09 0.95 0.93 0.16 0.85 0.82 0.46 0.88 0.78 0.27
RN50x64 0.97 0.96 0.09 0.96 0.95 0.12 0.96 0.94 0.13 0.90 0.84 0.23
ViT-B/32 0.90 0.88 0.32 0.87 0.84 0.37 0.91 0.88 0.27 0.90 0.84 0.23
ViT-B/16 0.91 0.89 0.25 0.89 0.85 0.31 0.88 0.85 0.34 0.93 0.88 0.16
ViT-L/14 0.91 0.88 0.27 0.88 0.85 0.38 0.86 0.82 0.35 0.92 0.86 0.20
DINO ViT-S/16 0.90 0.85 0.23 0.89 0.83 0.28 0.82 0.76 0.45 0.87 0.82 0.36
ViT-S/8 0.85 0.77 0.31 0.85 0.77 0.32 0.76 0.67 0.48 0.82 0.78 0.50
ViT-B/16 0.88 0.82 0.26 0.88 0.82 0.28 0.87 0.82 0.33 0.94 0.92 0.23
ViT-B/8 0.93 0.88 0.18 0.92 0.87 0.20 0.78 0.71 0.48 0.91 0.88 0.27
RN50 0.90 0.87 0.34 0.84 0.82 0.57 0.92 0.90 0.26 0.71 0.68 0.76
DINOv2 ViT-S/14 0.89 0.83 0.25 0.85 0.78 0.35 0.77 0.71 0.63 0.79 0.70 0.53
ViT-B/14 0.86 0.79 0.35 0.85 0.78 0.39 0.85 0.79 0.40 0.82 0.75 0.46
ViT-L/14 0.84 0.78 0.42 0.85 0.79 0.39 0.84 0.78 0.48 0.85 0.79 0.39
ViT-G/14 0.87 0.81 0.35 0.88 0.83 0.31 0.89 0.86 0.31 0.89 0.86 0.31
Grounding DINO SwinB 0.96 0.95 0.13 0.97 0.96 0.13 0.98 0.97 0.10 0.99 0.99 0.02
SwinT 0.96 0.95 0.13 0.96 0.95 0.14 0.99 0.99 0.04 0.98 0.98 0.08
Table 12: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Fog and ACDC Night [66] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.89 0.86 0.36 0.90 0.87 0.32 0.93 0.92 0.24 0.94 0.90 0.21
RN101 0.87 0.84 0.42 0.87 0.84 0.39 0.91 0.89 0.30 0.91 0.87 0.25
RN50x4 0.89 0.86 0.36 0.90 0.88 0.30 0.93 0.92 0.22 0.93 0.89 0.21
RN50x16 0.89 0.86 0.31 0.90 0.87 0.28 0.91 0.89 0.28 0.93 0.89 0.19
RN50x64 0.95 0.94 0.17 0.96 0.96 0.14 0.92 0.90 0.26 0.92 0.87 0.22
ViT-B/32 0.85 0.83 0.41 0.87 0.84 0.39 0.92 0.90 0.26 0.95 0.92 0.18
ViT-B/16 0.84 0.79 0.44 0.86 0.81 0.38 0.91 0.89 0.27 0.90 0.85 0.25
ViT-L/14 0.84 0.78 0.41 0.87 0.81 0.35 0.86 0.83 0.42 0.92 0.87 0.23
DINO ViT-S/16 0.78 0.73 0.64 0.80 0.76 0.60 0.82 0.80 0.61 0.84 0.84 0.52
ViT-S/8 0.80 0.77 0.62 0.82 0.80 0.54 0.83 0.83 0.55 0.86 0.87 0.56
ViT-B/16 0.81 0.77 0.56 0.84 0.80 0.53 0.85 0.84 0.53 0.81 0.83 0.68
ViT-B/8 0.82 0.80 0.51 0.85 0.82 0.47 0.84 0.84 0.52 0.84 0.86 0.58
RN50 0.66 0.62 0.78 0.59 0.57 0.88 0.70 0.67 0.81 0.55 0.54 0.91
DINOv2 ViT-S/14 0.75 0.71 0.69 0.78 0.74 0.70 0.79 0.78 0.66 0.84 0.82 0.62
ViT-B/14 0.70 0.67 0.74 0.71 0.68 0.71 0.71 0.70 0.78 0.73 0.69 0.75
ViT-L/14 0.63 0.61 0.80 0.67 0.65 0.79 0.67 0.66 0.83 0.75 0.72 0.67
ViT-G/14 0.64 0.61 0.77 0.71 0.67 0.69 0.72 0.69 0.71 0.78 0.73 0.58
Grounding DINO SwinB 0.75 0.70 0.63 0.72 0.66 0.67 0.77 0.73 0.60 0.89 0.85 0.35
SwinT 0.71 0.66 0.71 0.67 0.62 0.68 0.81 0.76 0.52 0.89 0.86 0.34
@@
CLIP RN50 0.98 0.97 0.08 0.98 0.97 0.08 0.99 0.99 0.05 0.98 0.96 0.08
RN101 0.96 0.94 0.14 0.96 0.95 0.12 0.98 0.97 0.09 0.98 0.97 0.06
RN50x4 0.98 0.97 0.08 0.98 0.98 0.06 0.99 0.99 0.07 0.99 0.99 0.04
RN50x16 0.98 0.97 0.07 0.98 0.97 0.06 0.99 0.99 0.04 0.98 0.96 0.05
RN50x64 0.99 0.98 0.05 0.99 0.99 0.04 0.99 0.98 0.05 0.99 0.98 0.03
ViT-B/32 0.96 0.96 0.12 0.97 0.96 0.11 0.99 0.99 0.05 0.99 0.99 0.03
ViT-B/16 0.95 0.94 0.16 0.95 0.94 0.15 0.99 0.99 0.05 0.99 0.98 0.05
ViT-L/14 0.93 0.90 0.20 0.94 0.92 0.15 0.97 0.97 0.11 0.98 0.97 0.05
DINO ViT-S/16 0.81 0.76 0.46 0.83 0.79 0.45 0.87 0.84 0.43 0.86 0.83 0.51
ViT-S/8 0.76 0.70 0.58 0.79 0.73 0.52 0.81 0.76 0.48 0.89 0.87 0.41
ViT-B/16 0.82 0.76 0.48 0.84 0.79 0.42 0.86 0.83 0.43 0.88 0.86 0.48
ViT-B/8 0.83 0.78 0.40 0.87 0.82 0.33 0.89 0.86 0.38 0.92 0.92 0.34
RN50 0.80 0.76 0.59 0.74 0.72 0.72 0.82 0.78 0.59 0.61 0.60 0.90
DINOv2 ViT-S/14 0.91 0.86 0.22 0.94 0.90 0.18 0.94 0.93 0.28 0.96 0.93 0.15
ViT-B/14 0.81 0.76 0.51 0.85 0.80 0.43 0.87 0.85 0.49 0.94 0.91 0.23
ViT-L/14 0.74 0.67 0.63 0.78 0.72 0.55 0.79 0.74 0.62 0.88 0.83 0.38
ViT-G/14 0.71 0.65 0.66 0.78 0.72 0.48 0.82 0.77 0.49 0.91 0.86 0.28
Grounding DINO SwinB 0.79 0.75 0.59 0.80 0.75 0.52 0.93 0.92 0.24 0.99 0.98 0.04
SwinT 0.76 0.71 0.59 0.75 0.70 0.59 0.94 0.92 0.24 0.98 0.97 0.07
Table 13: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Rain and ACDC Snow [66] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, OC-SVM, GMM, and NF methods. Best values in bold.

### v3

@@ D.2 Covariate Shift Experiments
CLIP RN50 0.94 0.91 0.17 0.94 0.91 0.17 0.92 0.89 0.23 0.93 0.90 0.27 0.92 0.84 0.19
RN101 0.96 0.94 0.15 0.96 0.94 0.16 0.92 0.89 0.25 0.87 0.83 0.39 0.91 0.84 0.22
RN50x4 0.98 0.98 0.07 0.98 0.98 0.07 0.94 0.93 0.19 0.87 0.83 0.43 0.83 0.75 0.34
RN50x16 0.97 0.96 0.09 0.97 0.96 0.10 0.95 0.93 0.16 0.85 0.82 0.46 0.88 0.78 0.27
RN50x64 0.97 0.96 0.09 0.97 0.96 0.09 0.96 0.95 0.12 0.96 0.94 0.13 0.90 0.84 0.23
ViT-B/32 0.90 0.88 0.32 0.90 0.88 0.32 0.87 0.84 0.37 0.91 0.88 0.27 0.90 0.84 0.23
ViT-B/16 0.91 0.89 0.25 0.91 0.89 0.26 0.89 0.85 0.31 0.88 0.85 0.34 0.93 0.88 0.16
ViT-L/14 0.91 0.88 0.27 0.91 0.88 0.27 0.88 0.85 0.38 0.86 0.82 0.35 0.92 0.86 0.20
DINO ViT-S/16 0.90 0.85 0.23 0.90 0.85 0.23 0.89 0.83 0.28 0.82 0.76 0.45 0.87 0.82 0.36
ViT-S/8 0.85 0.77 0.31 0.85 0.77 0.31 0.85 0.77 0.32 0.76 0.67 0.48 0.82 0.78 0.50
ViT-B/16 0.88 0.82 0.26 0.88 0.82 0.26 0.88 0.82 0.28 0.87 0.82 0.33 0.94 0.92 0.23
ViT-B/8 0.93 0.88 0.18 0.93 0.88 0.18 0.92 0.87 0.20 0.78 0.71 0.48 0.91 0.88 0.27
RN50 0.90 0.87 0.34 0.90 0.87 0.34 0.84 0.82 0.57 0.92 0.90 0.26 0.71 0.68 0.76
DINOv2 ViT-S/14 0.89 0.83 0.25 0.89 0.83 0.25 0.85 0.78 0.35 0.77 0.71 0.63 0.79 0.70 0.53
ViT-B/14 0.86 0.79 0.35 0.86 0.79 0.35 0.85 0.78 0.39 0.85 0.79 0.40 0.82 0.75 0.46
ViT-L/14 0.84 0.78 0.42 0.84 0.78 0.42 0.85 0.79 0.39 0.84 0.78 0.48 0.85 0.79 0.39
ViT-G/14 0.87 0.81 0.35 0.87 0.81 0.35 0.88 0.83 0.31 0.89 0.86 0.31 0.89 0.86 0.31
Grounding DINO SwinB 0.96 0.95 0.13 0.96 0.95 0.13 0.97 0.96 0.13 0.98 0.97 0.10 0.99 0.99 0.02
SwinT 0.96 0.95 0.13 0.96 0.95 0.13 0.96 0.95 0.14 0.99 0.99 0.04 0.98 0.98 0.08
Table 13: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Fog and ACDC Night [69] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.
@@
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50 0.89 0.86 0.36 0.89 0.86 0.37 0.90 0.87 0.32 0.93 0.92 0.24 0.94 0.90 0.21
RN101 0.87 0.84 0.42 0.87 0.84 0.42 0.87 0.84 0.39 0.91 0.89 0.30 0.91 0.87 0.25
RN50x4 0.89 0.86 0.36 0.89 0.86 0.36 0.90 0.88 0.30 0.93 0.92 0.22 0.93 0.89 0.21
RN50x16 0.89 0.86 0.31 0.89 0.86 0.31 0.90 0.87 0.28 0.91 0.89 0.28 0.93 0.89 0.19
RN50x64 0.95 0.94 0.17 0.95 0.94 0.17 0.96 0.96 0.14 0.92 0.90 0.26 0.92 0.87 0.22
ViT-B/32 0.85 0.83 0.41 0.85 0.83 0.41 0.87 0.84 0.39 0.92 0.90 0.26 0.95 0.92 0.18
ViT-B/16 0.84 0.79 0.44 0.84 0.79 0.44 0.86 0.81 0.38 0.91 0.89 0.27 0.90 0.85 0.25
ViT-L/14 0.84 0.78 0.41 0.84 0.78 0.42 0.87 0.81 0.35 0.86 0.83 0.42 0.92 0.87 0.23
DINO ViT-S/16 0.78 0.73 0.64 0.78 0.73 0.64 0.80 0.76 0.60 0.82 0.80 0.61 0.84 0.84 0.52
ViT-S/8 0.80 0.77 0.62 0.80 0.77 0.62 0.82 0.80 0.54 0.83 0.83 0.55 0.86 0.87 0.56
ViT-B/16 0.81 0.77 0.56 0.81 0.77 0.56 0.84 0.80 0.53 0.85 0.84 0.53 0.81 0.83 0.68
ViT-B/8 0.82 0.80 0.51 0.82 0.80 0.51 0.85 0.82 0.47 0.84 0.84 0.52 0.84 0.86 0.58
RN50 0.66 0.62 0.78 0.66 0.62 0.78 0.59 0.57 0.88 0.70 0.67 0.81 0.55 0.54 0.91
DINOv2 ViT-S/14 0.75 0.71 0.69 0.75 0.71 0.69 0.78 0.74 0.70 0.79 0.78 0.66 0.84 0.82 0.62
ViT-B/14 0.70 0.67 0.74 0.70 0.67 0.74 0.71 0.68 0.71 0.71 0.70 0.78 0.73 0.69 0.75
ViT-L/14 0.63 0.61 0.80 0.63 0.61 0.80 0.67 0.65 0.79 0.67 0.66 0.83 0.75 0.72 0.67
ViT-G/14 0.64 0.61 0.77 0.64 0.61 0.77 0.71 0.67 0.69 0.72 0.69 0.71 0.78 0.73 0.58
Grounding DINO SwinB 0.75 0.70 0.63 0.76 0.71 0.62 0.72 0.66 0.67 0.77 0.73 0.60 0.89 0.85 0.35
SwinT 0.71 0.66 0.71 0.71 0.66 0.70 0.67 0.62 0.68 0.81 0.76 0.52 0.89 0.86 0.34
@@
CLIP RN50 0.98 0.97 0.08 0.98 0.97 0.08 0.98 0.97 0.08 0.99 0.99 0.05 0.98 0.96 0.08
RN101 0.96 0.94 0.14 0.96 0.94 0.14 0.96 0.95 0.12 0.98 0.97 0.09 0.98 0.97 0.06
RN50x4 0.98 0.97 0.08 0.98 0.97 0.08 0.98 0.98 0.06 0.99 0.99 0.07 0.99 0.99 0.04
RN50x16 0.98 0.97 0.07 0.98 0.97 0.07 0.98 0.97 0.06 0.99 0.99 0.04 0.98 0.96 0.05
RN50x64 0.99 0.98 0.05 0.99 0.98 0.05 0.99 0.99 0.04 0.99 0.98 0.05 0.99 0.98 0.03
ViT-B/32 0.96 0.96 0.12 0.96 0.96 0.12 0.97 0.96 0.11 0.99 0.99 0.05 0.99 0.99 0.03
ViT-B/16 0.95 0.94 0.16 0.95 0.94 0.16 0.95 0.94 0.15 0.99 0.99 0.05 0.99 0.98 0.05
ViT-L/14 0.93 0.90 0.20 0.93 0.90 0.20 0.94 0.92 0.15 0.97 0.97 0.11 0.98 0.97 0.05
DINO ViT-S/16 0.81 0.76 0.46 0.81 0.76 0.46 0.83 0.79 0.45 0.87 0.84 0.43 0.86 0.83 0.51
ViT-S/8 0.76 0.70 0.58 0.76 0.70 0.58 0.79 0.73 0.52 0.81 0.76 0.48 0.89 0.87 0.41
ViT-B/16 0.82 0.76 0.48 0.82 0.76 0.48 0.84 0.79 0.42 0.86 0.83 0.43 0.88 0.86 0.48
ViT-B/8 0.83 0.78 0.40 0.83 0.78 0.40 0.87 0.82 0.33 0.89 0.86 0.38 0.92 0.92 0.34
RN50 0.80 0.76 0.59 0.80 0.76 0.59 0.74 0.72 0.72 0.82 0.78 0.59 0.61 0.60 0.90
DINOv2 ViT-S/14 0.91 0.86 0.22 0.91 0.86 0.22 0.94 0.90 0.18 0.94 0.93 0.28 0.96 0.93 0.15
ViT-B/14 0.81 0.76 0.51 0.81 0.76 0.51 0.85 0.80 0.43 0.87 0.85 0.49 0.94 0.91 0.23
ViT-L/14 0.74 0.67 0.63 0.74 0.67 0.63 0.78 0.72 0.55 0.79 0.74 0.62 0.88 0.83 0.38
ViT-G/14 0.71 0.65 0.66 0.71 0.65 0.66 0.78 0.72 0.48 0.82 0.77 0.49 0.91 0.86 0.28
Grounding DINO SwinB 0.79 0.75 0.59 0.79 0.75 0.58 0.80 0.75 0.52 0.93 0.92 0.24 0.99 0.98 0.04
SwinT 0.76 0.71 0.59 0.76 0.71 0.60 0.75 0.70 0.59 0.94 0.92 0.24 0.98 0.97 0.07
Table 14: VFM performance comparison across (i.e. ResNet, ViT, and Swin) architectures on ACDC Rain and ACDC Snow [69] datasets. AUROC↑, AUPR↑, and FPR95↓ metrics shown for APS, MFS, OC-SVM, GMM, and NF methods. Best values in bold.